OpenStudy (anonymous):

given that f(x) = x^2 - 81x and g(x) = x + 9 determine the simplified equation for the combined function (f(x)/g(x))

OpenStudy (alexandervonhumboldt2):

\[\frac{ x^2-81x }{ x+9 }\]

OpenStudy (alexandervonhumboldt2):

hmm

OpenStudy (nincompoop):

\(f(x) \div g(x)\rightarrow \dfrac{x^2-81x}{x+9} \)

OpenStudy (nincompoop):

you can try to factor the numerator and see if you can simplify

OpenStudy (alexandervonhumboldt2):

factor out the common term.

OpenStudy (anonymous):

do you know what x =/= (is not equal to)

OpenStudy (alexandervonhumboldt2):

hint: you cannot divide by 0. thus x+9 cannot be 0. solve for x to know what value cannot stand for x.

OpenStudy (alexandervonhumboldt2):

\[x+9\neq 0\]

OpenStudy (anonymous):

ah ok

OpenStudy (nincompoop):

do the quotient and reminder

OpenStudy (nincompoop):

remaindeR*

OpenStudy (nincompoop):

\((x-90)(x+9)+810\)

OpenStudy (alexandervonhumboldt2):

so we have this buddy divided by x+9.

OpenStudy (nincompoop):

yes, which still leads to the simplification by fraction if we want to look for value of x

OpenStudy (nincompoop):

\(x(x-81): x = 0; x=81 \)

OpenStudy (alexandervonhumboldt2):

hey gane :)

OpenStudy (nincompoop):

why is she saying not equal?

OpenStudy (alexandervonhumboldt2):

because there is a domain: x cannot be 9. otherwise we get infinity. :(

OpenStudy (alexandervonhumboldt2):

wopps i meant -9

OpenStudy (nincompoop):

oh

OpenStudy (alexandervonhumboldt2):

happens :)

OpenStudy (alexandervonhumboldt2):

it is wrong :(

OpenStudy (nincompoop):

conjugate?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (nincompoop):

I thought of that but I was lazy to try it out

OpenStudy (alexandervonhumboldt2):

\but wait it doesn't simplify :(

OpenStudy (nincompoop):

I mean if we are just looking for values of x, it is easy factor or complete the square alone can take of that but to evaluate, the function

OpenStudy (alexandervonhumboldt2):

maybe it is x^2-81, not x^2-81 xD .

OpenStudy (alexandervonhumboldt2):

i meant x^2-81x

OpenStudy (nincompoop):

it is x^2 - 81x no?

OpenStudy (alexandervonhumboldt2):

it is :(

OpenStudy (nincompoop):

complete the square for me

OpenStudy (nincompoop):

the numerator

OpenStudy (alexandervonhumboldt2):

not fun nin: \[(x-\frac{ 81 }{ 2 })^2+\frac{ 6564 }{ 4 }\]

OpenStudy (nincompoop):

lolllll

OpenStudy (alexandervonhumboldt2):

wat lol i broke my mind solving this buddy xD

OpenStudy (alexandervonhumboldt2):

i will never believe this will help us simplify out expression

OpenStudy (alexandervonhumboldt2):

GRRR